An electric motor with an efficiency 0.98 is used to lift a object with a weight of 50.0 N. How munch input power is needed to lift the object 12.0 m?

a. 600 J

b. 600 W

c. 612 J

d. 588 W

e. Not enough information given

A 25.0 lbm object is weighed on a planet that has a gravitational field of unknown strength. If the weight is determined through measurement to be 22.5 lbf, what is the acceleration due to gravity on this planet?

a. 1.11 ft/s^2

b. 35.8 ft/s^2

c. 0.900 ft/s^2

d. 29.0 ft/s^2

e. Not enough information given
Answer #1

In first part information is incomplete as power is asked but total time of lift is not given. Hence option e is correct

B) W = mg or w is directly proportional to g

Hence W1/W2= g1/g2

Or 25×32.174/22.5 = 32.174 / g2

So g2= 1.11 ft/s^2

Hence option a is correct